OpenStudy (anonymous):

Use the definitions and theorems of this section to evaluate and simplify the following expression. Be sure to express answers with positive exponents. (b^(3))^(2)

OpenStudy (zehanz):

Split it up: (b³)²=b³*b³ b³ = b*b*b So (b³)² = ...

OpenStudy (anonymous):

 is that the answer

OpenStudy (zehanz):

If this is it: I agree!